OpenStudy (anonymous):

GCF (8,36) GCF (50,75) What angle is formed by two right angles who have exactly one ray in common? Please help me?

OpenStudy (zzr0ck3r):

36=8*4+4 8=4*2+0 so gcf(8,36)=4 75=50*1+25 50=25*2+0 so gcd(50,75) = 25 do you see how I am doing this?

OpenStudy (anonymous):

Yes, thank you for showing me how to do that. I appreciate it!